# ACDiT: Interpolating Autoregressive Conditional Modeling and Diffusion Transformer

### v1

@@ Abstract
The recent surge of interest in comprehensive multimodal models has necessitated the unification of diverse modalities. However, the unification suffers from disparate methodologies. Continuous visual generation necessitates the full-sequence-attend diffusion-based approach, despite its divergence from the autoregressive modeling in the text domain. In this paper, we explore an interpolation between autoregressive and modeling to model visual information. At its core, we present ACDiT, an Autoregressive blockwise Conditional Diffusion Transformer, where the block size of diffusion, i.e., the size of autoregressive units, can be flexibly adjusted to interpolate between token-wise autoregression and full-sequence diffusion denoising. ACDiT is easy to implement, as simple as creating a Skip-Causal Attention Mask (SCAM) during training. During inference, the process iterates between diffusion denoising and autoregressive decoding that can make full use of KV-Cache. We show that ACDiT performs best among all autoregressive baselines under similar model scales on image and video generation tasks. We also demonstrate that benefiting from autoregressive modeling, pretrained ACDiT can be transferred in visual understanding tasks despite being trained with the diffusion objective. The analysis of the trade-off between autoregressive modeling and diffusion demonstrates the potential of ACDiT to be used in long-horizon visual generation tasks. These strengths make it promising as the backbone of future unified models.
@@ 1 Introduction
The concept of predicting the future has been a fundamental principle of Artificial Intelligence, ranging from the “next token prediction” objective of autoregressive language modeling [52, 6, 1] to predicting the next action in Reinforcement Learning [46, 59, 12, 53]. The recent renascence of the concept “World Model” also builds upon forming an internal state that is both capable of predicting the future and influenced by the prediction of the future [23]. The success of Large Language Models (LLMs) [67, 68, 6, 1, 32, 2, 78] also exemplify the power of such philosophy, demonstrating that complex abilities can arise from this simple objective.
@@
Existing works attempt to unify modalities by converting multimodal generation tasks into discrete token prediction tasks with vector quantization techniques [19, 65, 72, 66] and training on the mixed sequences with a next-token prediction objective. However, approaching the continuous distribution requires huge vocabulary sizes and a high utilization rate [81, 73], which is a complex objective. The information loss during vector quantization poses challenges for visual understanding tasks that require detailed information, such as the Optical Character Recognition task.
As a first step for exploring a unified framework, we propose ACDiT, an Autoregressive blockwise Conditional Diffusion Transformer that fuses the diffusion process with the autoregressive paradigm. At a high level, we extend the autoregressive units from the individual text token to blocks. The generation of each block can be formulated as a conditional diffusion process based on the previous block, where each block consists of visual patches of flexible size. As Fig. 1 shows, for image generation, a block can be defined as a small part of the image, while for video, a block can be a frame or multiple frames.
ACDiT is easy to implement, as simple as adding a Skip-Causal Attention Mask to the current DiT architecture during training. The inference process is formatted as an iteration between the conditional diffusion denoising process within a block, conditioned on the complete clean context, and autoregressive generation of a new block appended as the new context. In this way, KV-Cache can be used for faster inference. In general, ACDiT offers the following inherent advantages: (i) ACDiT simultaneously learns the causal interdependence across blocks with autoregressive modeling and the non-causal dependence within blocks with diffusion modeling. This serves as a versatile framework for expanding into unified multimodal and world models without conflict. (ii) ACDiT is endowed with clean continuous visual input, which can benefit visual understanding tasks in multimodal models. (iii) ACDiT makes full use of KV-Cache for flexible autoregressive generation in any length and potentially other latest long-context techniques in text for long video generation.
We validate ACDiT in both image and video generation tasks on ImageNet [57] and UCF101 [62], respectively. Experimental results demonstrate that ACDiT outperforms all autoregressive baselines of the same model scale and achieves visual quality comparable to full-sequence diffusion models, while exhibiting significantly higher inference speed when extended to long sequences. The codes and models are available at https://github.com/thunlp/ACDiT.
@@ 2 Related Work
Autoregressive Visual Generation. Autoregressive models have shown promising results in visual generation [13, 19, 66, 39, 72, 40]. The iGPT [13] first proposes autoregressively generating raw image pixels as a raster-scan sequence. VQGAN [19] improves the performance by training an autoregressive transformer on discrete tokens produced by VQVAE [69]. LlamaGen [64] enhances the image tokenizer and scales up the autoregressive transformers to 3.1B parameters, building on the latest Llama architecture [67, 68]. Also, several works demonstrate the potential of token-base visual generation in text-to-image tasks [80, 41, 64, 72]. Inspired by RQ-Transformer [36], VAR [66] proposes the next-scale prediction and obtain good improvement. In video generation, some works [29, 56] utilize sliding windows for progressive generation. Diffusion Forcing [8] and MAR [39] are two highly related works. Diffusion Forcing trains a causal autoregressive model to generate blocks without fully diffusing past ones and implement it on small RNN. MAR proposes the diffusion loss to learn the autoregressive conditional distribution on the head of the main Transformer with a small MLP network. Differently, ACDiT generates each block based on clear past and utilizes the full parameters of Transformer to denoise each block.
Unified Model for Understanding and Generation. Unified models for visual understanding and generation have recently garnered widespread attention. Some early efforts have aimed to align both the visual encoder and visual decoder with pre-trained Large Language Models [15, 75]. Some works utilize the discrete visual token to unify the image understanding and generation tasks, such as VILA-U [76], EMU3 [72], and Show-o [77]. Transfusion [82] tries the first attempt of joint training with language modeling loss and diffusion loss in a single transformer.
World Model. World models obtain great attention in various domains. Genie [7] introduces interactive game video simulation by learning latent actions on unlabeled video. Unisim [79] builds real-world interaction simulations with web video data. iVideoGPT [74] train an autoregressive transformer on a mixed sequence of tokens organized with visual observations, actions, and rewards. Despite the success, world models face challenges in leveraging the powerful generative capabilities of diffusion models due to their non-autoregressive nature.
@@
Figure 2: (a): for each noised block $n_{i}$, it can only attend previous clean latent $c_{0},c_{1},\dots,c_{i-1}$ and itself. Each clean block $c_{i}$ only attends the previous clean latent block. (b): ACDiT can effectively utilize the KC-Cache for autoregressive inference. (c): the 3D view of ACDiT, where $B,L,T$ denotes the block size, number of blocks, and denoising time steps, respectively. Darker color denotes higher noises.
@@ 3.1 Autoregressive Modeling
Autoregression asserts that the value at each timestep is contingent upon its preceding values. This principle is exemplified in autoregressive language models, which iteratively predict the probability distribution of subsequent tokens. Given a sequence of tokens ($x_{1},x_{2},\dots,x_{n}$), a salient characteristic of autoregression is that the prediction of $x_{i}$ is only dependent on its prefix ($x_{1},x_{2},\dots,x_{i-1}$). Upon determining $c_{i}$, it is concatenated with the preceding sequence, thereby forming the conditioning context ($x_{1},x_{2},\dots,x_{i}$) for predicting $x_{i+1}$. Therefore, the likelihood of sequence can be factorized as:
@@ 3.3 Desiderata
    The generation of future elements should be predicated on a precise representation of antecedent sequences. This is imperative because any ambiguity in the past inevitably complicates future predictions. This approach preserves the efficacy of autoregressive modeling and potentially facilitates the development of an “internal world model.” Furthermore, adherence to this principle enhances performance in discriminative tasks (e.g., visual understanding), as these tasks necessitate the input of all observable features into the model.
@@
Based on these desiderata, we analyze representative existing autoregressive diffusion methods. Diffusion-Forcing [8] proposes to use different-level random noise in the different positions of a sequence. Thus, in the inference process, denoising the subsequent positions from a clean past and be seen as a special case denoising different levels of noise. Their method does not meet the first desideratum. In the training process, the future is not predicted from the precise representation of the past. In MAR [39], the diffusion process is trained based on the latent in the last position, which does not satisfy the third desideratum. Moreover, the diffusion process sorely leverages the head part of the network, which conflicts with the second desideratum. In Transfusion [82], despite that the diffusion utilizes the full network parameters for both autoregressive modeling and diffusion, it does not utilize the “predicting the future” objective within the multimodal information. Moreover, when trained with multiple blocks of multimodal information, such as multiple images, the latter image will attend to the noise version of the former image. To comprise this deficiency, they use half of the noise schedule, i.e., limiting maximum noise steps to 500 instead of 1000, in 20% image captioning pairs, which is not an optimal strategy.
@@ 4.1 Framework
During the inference phase, each autoregressive step executes a conditional diffusion process for $n_{i}$ based on $\{c_{j}|j=0,...,i-1\}$’s KV-Cache. Upon finishing denoising, it is appended to the clean sequence as $c_{i}$ followed by the maximal noise version of the next block $n_{i+1}$. The key-value tensor will be computed for these two blocks, and the key-value tensor of the clean block $c_{i}$ will be kept in KV-Cache. All noise-corrupted version of $n_{i}$ is disregarded. The process is visualized in Figure 8. A three-dimensional view of our method is presented in Figure 2(c). By interpolating between full-sequence diffusion and autoregressive paradigms, ACDiT enjoys flexibility and expressivity, enabling it to generate video of any length utilizing the latest long-context techniques developed for language models.
@@ 4.2 Positional Encoding
In this formulation, each $\mathbf{R}^{d_{j}}_{\Theta_{j},m_{j}}$ represents a $d_{j}$-dimensional rotary matrix [^2] For a detailed explanation, please refer to Equation 15 in version 5 of the arXiv preprint of  [63]. with rotation angles $\Theta_{j}=\{\theta_{ji}=b_{j}^{-2(i-1)/d_{j}},i\in[1,2,\cdots,\frac{d_{j}}{2}]\}$. The base $b_{j}$ is empirically determined as $100\lceil\frac{8L_{j}}{100\pi}\rceil$, where $L_{j}$ denotes the maximum position index in data dimension $j$. This formulation ensures that the highest wavelength of RoPE is approximately eight times the maximum position, thereby mitigating rapid decay in long-term dependencies.
@@ 4.3 Efficiency Analysis and Block Size Choice
Figure 4: The model architecture of ACDiT. Both the clean latent and noise latent are input while only noise tokens are scale shift with conditioning information.
Table 1: Image generation results on ImageNet 256x256.
@@
ImageFolder [40] (arxiv’24) AR Disc. 362M 2.60 295.0 0.75 0.63
@@ 4.4 Model Architecture
ACDiT mainly inherits the main architecture of DiT. However, since we want to keep the architecture as simple and unified as possible, we use linear layers instead of convolution in the input layer and final layer. Besides, we replace the absolute position embedding and Layer Normalization with RoPE [63] and RMSNorm (Root Mean Square Layer Normalization) [67], respectively. We find that QK-norm is important to stabilizing the video generation training, thus we use QK-norms in all experiments. The additional conditional information timesteps and labels are injected into the model with adaLN-Zero only on the noise part. For both image and video generation, we follow DiT and leverage the pre-trained image VAE [33] from Stable Diffusion [54], whose downsample factor is 8. For image generation under B block size, we group square latent representation patches with $\sqrt{B}\times\sqrt{B}$ shape as a block.
Table 2: Configuration details of ACDiT.
Model #Layers Hidden Size MLP #Heads Params
ACDiT-B 12 768 3072 12 132M
ACDiT-L 24 1024 4096 16 460M
ACDiT-XL 28 1152 4608 18 677M
ACDiT-H 32 1280 5120 20 954M
@@ 5 Experiments
Figure 5: Sample images from ACDiT-H on ImageNet 256$\times$256.
Figure 6: Sample videos from ACDiT-XL trained on UCF-101.
@@ 5.1 Experimental Setup
Implementation details. In the image generation task, we set the patch size as 1 and the autoregressive unit block size as $256=16\times 16$. Therefore, for a $256\times 256\times 3$ image in $32\times 32\times 4$ latent shape, the total sequence length and autoregressive length are 1024 and 4, respectively. We explore 4 different model sizes, as shown in Table 2. ACDiT-B is used for design verification and analyze. ACDiT is trained on ImageNet for 1.2M iterations with a batch size of 1024. We use the AdamW optimizer [42] and WSD (Warmup Steady Decay) learning rate scheduling [31] with the peak learning rate 3e-4 and no weight decay. The learning rate begins to decay in the last 15% training iteration. Following the common training recipe of generative models, we keep an exponential moving average (EMA) of the ACDiT weights during training using a decay rate of 0.9999. We sample images with DPM-Solver [43] for 25 steps within each block and use classifier-free guidance [27] with a guidance scale of 1.5. In video generation, we sample 16 frames from each video and set the patch size as 2 and the block size as $1024=256\times 4$. For a $16\times 256\times 256\times 3$ video in $16\times 32\times 32\times 4$ latent shape, the sequence length of each frame is 256 and the total sequence length is 4096, with 4 frames grouped into one block. We train ACDiT on UCF-101 for 400K iterations with a batch size of 96. The classifier-free guidance scale is 2.5. Other training configs are the same as image training. All models are implemented with PyTorch [48] and trained on NVIDIA H100 GPUs. Specifically, we use FlexAttention [^3] https://pytorch.org/blog/flexattention to implement the SCAM for both customization and efficiency.
@@
(a) Scaling performance of ACDiT.
(b) Ablation study for ROPE-ND.
(c) FID curve of last 30% training.

### v2

@@ Abstract
We present ACDiT, a novel Autoregressive blockwise Conditional Diffusion Transformer, that innovatively combines autoregressive and diffusion paradigms for modeling continuous visual information. By introducing a block-wise autoregressive unit, ACDiT offers a flexible interpolation between token-wise autoregression and full-sequence diffusion, bypassing the limitations of discrete tokenization. The generation of each block is formulated as a conditional diffusion process, conditioned on prior blocks. ACDiT is easy to implement, as simple as creating a Skip-Causal Attention Mask (SCAM) on standard diffusion transformer during training. During inference, the process iterates between diffusion denoising and autoregressive decoding that can make full use of KV-Cache. We show that ACDiT performs best among all autoregressive baselines under similar model scales on image and video generation tasks. We also demonstrate that benefiting from autoregressive modeling, pretrained ACDiT can be transferred in visual understanding tasks despite being trained with the diffusion objective. The analysis of the trade-off between autoregressive modeling and diffusion demonstrates the potential of ACDiT to be used in long-horizon visual generation tasks. We hope that ACDiT offers a novel perspective on visual autoregressive generation and unlocks new avenues for unified models.
@@ 1 Introduction
The concept of predicting the future has been a fundamental principle of Artificial Intelligence, ranging from the “next token prediction” objective of autoregressive language modeling [55, 6, 1] to predicting the next action in Reinforcement Learning [48, 62, 12, 57]. The success of Large Language Models (LLMs) [70, 71, 6, 1, 34, 2, 80] exemplify the power of such philosophy, demonstrating that complex abilities can arise from this simple objective.
@@
Existing works attempt to build the visual autoregressive model by converting visual generation tasks into discrete token prediction tasks with vector quantization techniques [21, 68, 75, 69] and training with a next-token prediction objective. However, approaching the continuous distribution requires huge vocabulary sizes and a high utilization rate [85, 76], which is a complex objective.
In this paper, we propose ACDiT, an Autoregressive blockwise Conditional Diffusion Transformer that fuses the diffusion process with the autoregressive paradigm. Instead of relying on discrete tokens, we demonstrate that visual autoregressive generation can also be achieved using continuous visual features. At a high level, we extend the autoregressive units from the individual discrete token to blocks, where each block consists of visual patches of flexible size. The generation of each block can be formulated as a conditional diffusion process based on the previous block. As Fig. 1 shows, for image generation, a block represents a small region of the image, whereas in video generation, a block can correspond to a single frame or multiple frames.
ACDiT is easy to implement, as simple as adding a Skip-Causal Attention Mask to the current diffusion transformer [52] architecture during training. The inference process is formatted as an iteration between the conditional diffusion denoising process within a block, conditioned on the complete clean context, and autoregressive generation of a new block appended as the new context. In this way, KV-Cache can be used for faster inference. In general, ACDiT offers the following inherent advantages: (i) ACDiT simultaneously learns the causal interdependence across blocks with autoregressive modeling and the non-causal dependence within blocks with diffusion modeling. (ii) ACDiT is endowed with clean continuous visual input, which eliminates the requirement for vector quantization and improves the transfering to visual understanding tasks after being trained with generative objectives. (iii) ACDiT makes full use of KV-Cache for flexible autoregressive generation in any length and can potentially benefit from other latest long-context techniques in text for long video generation.
We validate ACDiT in both image and video generation tasks on ImageNet [60] and UCF101 [65], respectively. Experimental results demonstrate that ACDiT outperforms all autoregressive baselines of the same model scale and achieves visual quality comparable to full-sequence diffusion models while exhibiting significantly higher inference speed when extended to long sequences. We hope ACDiT presents a fresh perspective on visual autoregressive generation and paves the way for exploring unified multimodal and world models. The codes and models are available at https://github.com/thunlp/ACDiT.
@@ 2 Related Work
Autoregressive Generation on Discrete Tokens. The iGPT [13] first proposes autoregressively generating raw image pixels as a raster-scan sequence. Most visual autoregressive models [56, 75, 83, 15, 16] follow VQGAN [21], which pioneers this direction by training an autoregressive transformer on discrete tokens produced by VQVAE [72]. LlamaGen [67] enhances the image tokenizer and scales up the autoregressive transformers building on the latest Llama architecture [70]. Inspired by RQ-Transformer [38], VAR [69] proposes the next-scale prediction and obtain good improvement. Some subsequent works [69, 43, 82] extend next-scale prediction to text-to-image generation.
Autoregressive Generation on Continous Features. Some recent works make some early exploration on achieving visual autoregressive generation without discrete tokens and combining the advantage of diffusion and autoregressive. Diffusion Forcing [8] trains a causal autoregressive model to generate blocks without fully diffusing past ones and implement it on small RNN. MAR [41] proposes the diffusion loss to learn the autoregressive conditional distribution on the head of the main Transformer with a small MLP network. Transfusion [87] and Monoformer [86] conduct joint training with language modeling loss and diffusion loss in a single transformer, while within image generation, their design is identical to the standard diffusion model. Dart [25] proposes to autoregressively denoised the complete image within in a Transformer. Different from these work, ACDiT redefine the autoregressive unit and generates each block based on clear past, providing a flexible framework for both image and video generation.
@@
Figure 2: (a): for each noised block $n_{i}$, it can only attend previous clean latent $c_{0},c_{1},\dots,c_{i-1}$ and itself. Each clean block $c_{i}$ only attends the previous clean latent block. (b): ACDiT can effectively utilize the KV-Cache for autoregressive inference. (c): the 3D view of ACDiT, where $B,L,T$ denotes the block size, number of blocks, and denoising time steps, respectively. The darker color denotes higher noises.
@@ 3.1 Autoregressive Modeling
Autoregression asserts that the value at each timestep is contingent upon its preceding values. This principle is exemplified in autoregressive language models, which iteratively predict the probability distribution of subsequent tokens. Given a sequence of tokens ($x_{1},x_{2},\dots,x_{n}$), a salient characteristic of autoregression is that the prediction of $x_{i}$ is only dependent on its prefix ($x_{1},x_{2},\dots,x_{i-1}$). Upon determining $x_{i}$, it is concatenated with the preceding sequence, thereby forming the conditioning context ($x_{1},x_{2},\dots,x_{i}$) for predicting $x_{i+1}$. Therefore, the likelihood of sequence can be factorized as:
@@ 3.3 Desiderata
    The generation of future elements should be predicated on a precise representation of antecedent sequences. This is imperative because any ambiguity in the past inevitably complicates future predictions. This approach preserves the efficacy of autoregressive modeling and potentially facilitates the development of the world model. Furthermore, adherence to this principle enhances performance in discriminative tasks (e.g., visual understanding), as these tasks necessitate the input of all observable features into the model.
@@
Based on these desiderata, we analyze representative existing autoregressive diffusion methods. Diffusion-Forcing [8] proposes to use different-level random noise in the different positions of a sequence. Thus, in the inference process, denoising the subsequent positions from a clean past can be seen as a special case denoising different levels of noise. Their method does not meet the first desideratum. In the training process, the future is not predicted from the precise representation of the past. In MAR [41], the diffusion process is trained based on the latent in the last position, which does not satisfy the third desideratum. Moreover, the diffusion process sorely leverages the head part of the network, which conflicts with the second desideratum. In Transfusion [87] and Monoformer [86], despite that the diffusion utilizes the full network parameters for both autoregressive modeling and diffusion, it does not utilize the “predicting the future” objective within the visual information. Moreover, when trained with multiple visual input, such as multiple images, the latter image will attend to the noise version of the former image. To comprise this deficiency, they half the noise schedule, i.e., limiting maximum noise steps to 500 from 1000, in 20% image captioning pairs, which is not an optimal strategy.
@@ 4.1 Framework
During the inference phase, each autoregressive step executes a conditional diffusion process for $n_{i}$ based on $\{c_{j}|j=0,...,i-1\}$’s KV-Cache. Upon finishing denoising, it is appended to the clean sequence as $c_{i}$ followed by the maximal noise version of the next block $n_{i+1}$. The key-value tensor will be computed for these two blocks, and the key-value tensor of the clean block $c_{i}$ will be kept in KV-Cache. All noise-corrupted version of $n_{i}$ is disregarded. The process is visualized in Figure 2(b). A three-dimensional view of our method is presented in Figure 2(c). By interpolating between full-sequence diffusion and autoregressive paradigms, ACDiT enjoys flexibility and expressivity, enabling it to generate video of any length utilizing the latest long-context techniques developed for language models.
@@ 4.2 Positional Encoding
In this formulation, each $\mathbf{R}^{d_{j}}_{\Theta_{j},m_{j}}$ represents a $d_{j}$-dimensional rotary matrix [^2] See a detailed explanation in the Equation 15 in  [66]. with rotation angles $\Theta_{j}=\{\theta_{ji}=b_{j}^{-2(i-1)/d_{j}},i\in[1,2,\cdots,\frac{d_{j}}{2}]\}$. The base $b_{j}$ is empirically determined as $100\lceil\frac{8L_{j}}{100\pi}\rceil$, where $L_{j}$ denotes the maximum position index in data dimension $j$. This formulation ensures that the highest wavelength of RoPE is approximately eight times the maximum position, thereby mitigating rapid decay in long-term dependencies.
@@ 4.3 Efficiency Analysis and Block Size Choice
Figure 4: The model architecture of ACDiT, mainly inherited from standard DiT architecture. Both the clean latent and noise latent are input while only noise tokens are scale shift with conditioning information. The SCAM attention mask pattern will be applied in the attention block.
Model #Layers Hidden Size MLP #Heads Params
ACDiT-B 12 768 3072 12 132M
ACDiT-L 24 1024 4096 16 460M
ACDiT-XL 28 1152 4608 18 677M
ACDiT-H 32 1280 5120 20 954M
Table 1: Configuration details of ACDiT.
@@
ImageFolder [42] (ICLR’25) AR Disc. 362M 2.60 295.0 0.75 0.63
LlamaGen+NPP [50] (arxiv’24) AR Cont. 1.4B 2.55 282.0 0.84 0.56
@@
MonoFormer [86] (arxiv’24) AR + Diff Cont. - 1.1B 2.57 272.6 0.84 0.56
Dart [26] (ICLR’25) AR + Diff Cont. 820M 3.82 263.8 - -
@@ 4.4 Model Architecture
As shown in Figure 4, ACDiT mainly inherits the main architecture of DiT. We replace the original bidirectional attention with the proposed SCAM attention pattern to process the clean and noise latent. Since we want to keep the architecture as simple and unified as possible, we use linear layers instead of convolution in the input layer and final layer. Besides, we replace the absolute position embedding and Layer Normalization with RoPE [66] and RMSNorm (Root Mean Square Layer Normalization) [70], respectively. We find that QK-norm is important to stabilizing the video generation training, thus we use QK-norms in all experiments. The additional conditional information timesteps and labels are injected into the model with adaLN-Zero only on the noise part. For both image and video generation, we follow DiT and leverage the pre-trained image VAE [35] from Stable Diffusion [58], whose downsample factor is 8. For image generation under B block size, we group square latent representation patches with $\sqrt{B}\times\sqrt{B}$ shape as a block.
@@ 5.1 Experimental Setup
Implementation details. In the image generation task, we set the patch size as 1 and the autoregressive unit block size as $256=16\times 16$. Therefore, for a $256\times 256\times 3$ image in $32\times 32\times 4$ latent shape, the total sequence length and autoregressive length are 1024 and 4, respectively. We explore 4 different model sizes, as shown in Table 1. ACDiT-B is used for design verification and analyze. ACDiT is trained on ImageNet for 1.2M iterations with a batch size of 1024. We use the AdamW optimizer [44] and WSD (Warmup Steady Decay) learning rate scheduling [33] with the peak learning rate 3e-4 and no weight decay. The learning rate begins to decay in the last 15% training iteration. Following the common training recipe of generative models, we keep an exponential moving average (EMA) of the ACDiT weights during training using a decay rate of 0.9999. We sample images with DPM-Solver [45] for 25 steps within each block and use classifier-free guidance [30] with a guidance scale of 1.5. In video generation, we sample 16 frames from each video and set the patch size as 2 and the block size as $1024=256\times 4$. For a $16\times 256\times 256\times 3$ video in $16\times 32\times 32\times 4$ latent shape, the sequence length of each frame is 256 and the total sequence length is 4096, with 4 frames grouped into one block. We train ACDiT on UCF-101 for 400K iterations with a batch size of 96. The classifier-free guidance scale is 2.5. Other training configs are the same as image training. All models are implemented with PyTorch [51]. Specifically, we use FlexAttention [^3] https://pytorch.org/blog/flexattention to implement the SCAM for both customization and efficiency. We use 256 and 96 NVIDIA H100 GPUs to train the image and video generation model, respectively.
(a) Scaling performance of ACDiT.
(b) Ablation study for ROPE-ND.
(c) FID curve of last 30% training.
@@
Figure 8: Sample images from ACDiT-H on ImageNet 256$\times$256.
Figure 9: Sample videos from ACDiT-XL trained on UCF-101.